# Differential Privacy Analysis of Decentralized Gossip Averaging under Varying Threat Models

### v1

@@ Appendix A Proof of Lemma 6
$=H_{\alpha}\left(\Sigma_{r}^{-1}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}+\widetilde{Z}\,||\,\widetilde{Z}\right),$
@@
$H_{{\rm e}\hskip 1.0pt^{\varepsilon}}\left(\mathcal{M}(D)||\mathcal{M}(D^{\prime})\right)=H_{{\rm e}\hskip 1.0pt^{\varepsilon}}\left(\Sigma_{r}^{-1}U_{r}^{\top}\big{(}f(D)-f(D^{\prime})\big{)}+\widetilde{Z}\,||\,\widetilde{Z}\right),$
@@ Appendix B Proof of Lemma 1
Lemma 1.
Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}-{x^{\prime}}_{t}^{j}\right|\leq 1$ (this without loss of generality). Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of node $i$, where $(\varepsilon,\delta)$ is bounded by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T})$, and
$\widetilde{x}_{T}=\begin{bmatrix}e_{i}^{\top}We_{j}\\ e_{i}^{\top}(W^{2}+W)e_{j}\\ \vdots\\ e_{i}^{\top}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{i}^{\top}W&0&0&\dots&0\\ e_{i}^{\top}W^{2}&e_{i}^{\top}W&0&\dots&0\\ e_{i}^{\top}W^{3}&e_{i}^{\top}W^{2}&e_{i}^{\top}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{i}^{\top}W^{T}&e_{i}^{\top}W^{T-1}&\dots&e_{i}^{\top}W^{2}&e_{i}^{\top}W\end{bmatrix}$
@@
We now need to bound the hockey-stick divergence
$H_{\alpha}\left(\mathrm{View}_{\mathcal{M}(D)}(\{i\})\,||\,\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\}))\right).$
Recall that from the recursion (3.1) it follows that
$\mathrm{View}_{\mathcal{M}(D)}(\{i\})=\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\end{bmatrix},$
where
$y_{1}$ $=e_{i}^{\top}(Wx_{0}+Wu_{0})$
$y_{2}$ $=e_{i}^{\top}(W^{2}x_{0}+Wx_{1}+W^{2}u_{0}+Wu_{1})$
$\vdots$
$y_{T}$ $=e_{i}^{\top}(W^{T}x_{0}+\cdots+Wx_{T-1}+W^{T}u_{0}+\cdots+Wu_{T-1}).$
I.e.,
$\mathrm{View}_{\mathcal{M}(D)}(\{i\})=H_{T}\widehat{x}_{T}+H_{T}\widetilde{u}_{T},$
where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T\cdot n})$ and
$\widehat{x}_{T}=\begin{bmatrix}x_{0}\\ x_{1}\\ \vdots\\ x_{T-1}\end{bmatrix}.$
similarly,
$\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\})=H_{T}\widehat{x}_{T}^{\prime}+H_{T}\widetilde{u}_{T},$
where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T\cdot n})$ and
$\widehat{x}_{T}^{\prime}=\begin{bmatrix}x_{0}^{\prime}\\ x_{1}^{\prime}\\ \vdots\\ x_{T-1}^{\prime}\end{bmatrix}$
for some $\widehat{x}_{T}^{\prime}\in\mathbb{R}^{T\cdot n}$ such that
$\widehat{x}_{T}-\widehat{x}_{T}^{\prime}=\begin{bmatrix}e_{j}\Delta_{1}\\ e_{j}\Delta_{2}\\ \vdots\\ e_{j}\Delta_{n}\end{bmatrix},$
where $e_{j}$ denotes the $j$th standard basis vector in $\mathbb{R}^{n}$ and $\left|\Delta_{i}\right|\leq 1$ for all $i\in[T]$.
@@
$=$ $H_{\alpha}\left(V_{r}^{\top}\big{(}\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big{)}+\widehat{u}\,||\,\widehat{u}\right)$
where $\widehat{u}\sim\mathcal{N}(0,\sigma^{2}I_{T})$.
Denoting $\Delta x:=\widehat{x}_{T}-\widehat{x}_{T}^{\prime}$, we see that $-1\leq(\Delta x)_{i}\leq 1$ for all $i\in[T]$. For an upper bound of the sensitivity, we need to find the maximum of $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ when $\Delta x$ is in the convex and compact domain $[0,1]^{T}$. We see that maximizing $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ is equivalent to finding the maximum of the quadratic form $(\Delta x)^{\top}P\Delta x$ in the convex and compact domain $[0,1]^{T}$ where the projector matrix $P=V_{r}V_{r}^{\top}$ is positive semi-definite and a projection on to the row space of $H_{T}$. Thus, the optimum is found from somewhere at the boundary, i.e., it holds that $\left|(\Delta x)_{i}\right|=1$ for all $i\in[T]$. Moreover, since the rows of $H_{T}$ are vectors with non-negative entries, $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ is maximized when $\Delta x=\begin{bmatrix}1&\ldots&1\end{bmatrix}^{\top}$. ∎
Appendix C Proof of Thm. 3
Theorem 1.
Consider the neighboring datasets $D$ and $D^{\prime}$ that change at the data of node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}(\theta)-{x^{\prime}}_{t}^{j}(\theta)\right|\leq 1$ for any auxiliary variable $\theta$. Denote
$\widetilde{x}_{T}=\begin{bmatrix}S_{j}We_{j}\\ S_{j}(W^{2}+W)e_{j}\\ \vdots\\ S_{j}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}S_{j}W&0&0&\dots&0\\ S_{j}W^{2}&S_{j}W&0&\dots&0\\ S_{j}W^{3}&S_{j}W^{2}&S_{j}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ S_{j}W^{T}&S_{j}W^{T-1}&\dots&S_{j}W^{2}&S_{j}W\end{bmatrix}$
Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of any other node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
@@
Suppose $D\simeq_{j}D^{\prime}$. Denote $S_{j}=S(\bar{N}_{j})$. We need to carry out the analysis for the view $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})=\begin{bmatrix}y_{1}&\ldots&y_{T}\end{bmatrix}^{\top},$ where
$y_{1}$ $=S_{j}\big{(}W(x_{0}+u_{0})\big{)}$
$y_{2}$ $=S_{j}\big{(}(W^{2}(x_{0}+u_{0})+W(x_{1}+u_{1})\big{)}$
$\vdots$
$y_{T}$ $=S_{j}\big{(}W^{T}(x_{0}+u_{0})+\cdots+W(x_{T-1}+u_{T-1})\big{)},$
where $[x_{i}]_{k}$ depends on the state $[\theta_{i}]_{k}$ (or could depend also on $[\theta_{i-1}]_{k},[\theta_{i-2}]_{k},\ldots$).
The only source of randomness in $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})$ is that of $u_{0},\ldots,u_{T-1}$. To carry out the DP analysis, we "roll out" by starting the integration of the hockey-stick divergence from the last noise vector $u_{T-1}$. This is similar to the composition analysis using RDP (Mironov,, 2017) and using dominating pairs of distributions (Zhu et al.,, 2022). In particular, the proof is similar to the proof of (Thm. 10, Zhu et al.,, 2022).
The general case can be illustrated using with the case $T=2$. When $T=2$, we need to analyze the mechanism
$\mathcal{M}(D)=\begin{bmatrix}S_{j}\big{(}W(x_{0}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(x_{0}+u_{0})+W(x_{1}+u_{1})\big{)}\end{bmatrix}=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},$
where $[x_{1}]_{k}$ depends on the state $[\theta_{1}]_{k}$ for all $k\in[n]$.
The essential observation here is that when $D\simeq_{j}D^{\prime}$, the the states of the nodes outside of the closed neighborhood $\bar{N}_{j}$ are obtained by applying the same post-processing to the views $\mathrm{View}_{\mathcal{M}(D)}(\bar{N}_{j})$ and $\mathrm{View}_{\mathcal{M}(D^{\prime})}(\bar{N}_{j})$. More specifically, given the view of the nodes inside the closed neighborhood $\bar{N}_{j}$, in case the output of the first component is $y_{1}$ and when only considering the randomness of the vector $W(x_{0}+u_{0})$ in the nodes outside of $\bar{N}_{j}$, the vectors $S_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $S_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ have the same distribution, since they consist of linear combinations of the entries of $y_{1}$ and of entries of $\theta_{1}$ outside of $\bar{N}_{j}$ which are distributed similarly in case of $D$ and $D^{\prime}$.
Therefore, when starting to integrate with respect to the last source of randomness, $u_{1}$, since $y_{1}$ is observed, we can fix both $S_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $S_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ to same value, and we can also fix all the terms in $Wx_{1}$ that are evaluated using the states of the nodes outside of $\bar{N}_{j}$. What remains in $Wx_{1}$ are the terms evaluated using $y_{1}$, and thus we can also carry out an adaptive analysis. We then roll out the last term to become state-independent, i.e., we have that for all $\alpha\geq 0$,
$H_{\alpha}\big{(}\mathcal{M}(D)||\mathcal{M}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)},$
where
$\mathcal{M}_{1}(D)=\begin{bmatrix}S_{j}\big{(}W(x_{0}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(x_{0}+u_{0})+W(e_{j}+u_{1})\big{)}\end{bmatrix}$
and
$\mathcal{M}_{1}(D^{\prime})=\begin{bmatrix}S_{j}\big{(}W(x_{0}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(x_{0}+u_{0})+Wu_{1}\big{)}\end{bmatrix}.$
Switching then the order of integration, we integrate out $u_{0}$ and, by the reasoning of the proof of Lemma 1, we have that
$H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{2}(D)||\mathcal{M}_{2}(D^{\prime})\big{)},$
where
$\mathcal{M}_{2}(D)=\begin{bmatrix}S_{j}\big{(}W(e_{j}+u_{0})\big{)}\\ S_{j}\big{(}W^{2}(e_{j}+u_{0})+W(e_{j}+u_{1})\big{)}\end{bmatrix}=\begin{bmatrix}S_{j}We_{j}\\ S_{j}\big{(}W^{2}+W\big{)}e_{j}\end{bmatrix}+\begin{bmatrix}S_{j}W&0\\ S_{j}W^{2}&S_{j}W\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}$
and
$\mathcal{M}_{2}(D^{\prime})=\begin{bmatrix}S_{j}\big{(}Wu_{0}\big{)}\\ S_{j}\big{(}W^{2}u_{0}+Wu_{1}\big{)}\end{bmatrix}=\begin{bmatrix}S_{j}W&0\\ S_{j}W^{2}&S_{j}W\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}.$
In case $T=3$, we can use the same reasoning, and similarly, given the observations $y_{1}$ and $y_{2}$, we can freeze the terms $S_{j}\big{(}W^{3}(x_{0}+u_{0})$ and $S_{j}W^{2}(x_{1}+u_{1})$ when integrating w.r.t. $u_{2}$. The general case also follows from this reasoning.
Appendix D Proof of Thm. 4
Theorem 1.
Consider the gossip averaging (3.5) corresponding to gossip averaging without secure summation. Denote
$\widetilde{x}_{T}=\begin{bmatrix}1\\ e_{j}^{\top}(W+I)e_{j}\\ \vdots\\ e_{j}^{\top}(W^{T-1}+\cdots+W+I)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{j}^{\top}&0&0&\dots&0\\ e_{j}^{\top}W&e_{j}^{\top}&0&\dots&0\\ e_{j}^{\top}W^{2}&e_{j}^{\top}W&e_{j}^{\top}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{j}^{\top}W^{T-1}&e_{j}^{\top}W^{T-2}&\dots&e_{j}^{\top}W&e_{j}^{\top}\end{bmatrix}$
Then, the node $i$ is $(\varepsilon,\delta)$-DP from the point-of-view of any node $i$, where $(\varepsilon,\delta)$ is the DP guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert H_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ and noise scale $\sigma$.
@@
The proof goes similarly as the proof of Thm. 3 and can be illustrated using the case $T=2$.
Given the view of the node $j$, in case the output of the first component is $y_{1}$ and when only considering the randomness of the vector $W(x_{0}+u_{0})$ in the rest of the nodes, the vectors $e_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $e_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ have the same distribution.
Therefore, when starting to integrate w.r.t. to the last source of randomness, $u_{1}$, since $y_{1}$ is observed, we can fix both $e_{j}\big{(}W^{2}(x_{0}+u_{0})$ and $e_{j}\big{(}W^{2}(x_{0}^{\prime}+u_{0})$ to same value, and roll out the last term to become state-independent, i.e., we have that for all $\alpha\geq 0$,
$H_{\alpha}\big{(}\mathcal{M}(D)||\mathcal{M}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)},$
where
$\mathcal{M}_{1}(D)=\begin{bmatrix}e_{j}^{\top}(x_{0}+u_{0})\\ e_{j}^{\top}\big{(}W(x_{0}+u_{0})+e_{j}+u_{1}\big{)}\end{bmatrix}$
and
$\mathcal{M}_{1}(D^{\prime})=\begin{bmatrix}e_{j}^{\top}(x_{0}+u_{0})\\ e_{j}^{\top}\big{(}W(x_{0}+u_{0})+u_{1}\big{)}\end{bmatrix}$
Switching then the order of integration, we integrate out $u_{0}$ and, by the reasoning of the proof of Lemma 1, we have that
$H_{\alpha}\big{(}\mathcal{M}_{1}(D)||\mathcal{M}_{1}(D^{\prime})\big{)}\leq H_{\alpha}\big{(}\mathcal{M}_{2}(D)||\mathcal{M}_{2}(D^{\prime})\big{)},$
where
$\mathcal{M}_{2}(D)=\begin{bmatrix}e_{j}^{\top}(e_{j}+u_{0})\\ e_{j}^{\top}\big{(}W(e_{j}+u_{0})+e_{j}+u_{1}\big{)}\end{bmatrix}=\begin{bmatrix}1\\ e_{j}^{\top}\big{(}W+I\big{)}e_{j}\end{bmatrix}+\begin{bmatrix}e_{j}^{\top}&0\\ e_{j}^{\top}W&e_{j}\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}$
and
$\mathcal{M}_{2}(D^{\prime})=\begin{bmatrix}e_{j}^{\top}Wu_{0}\\ e_{j}^{\top}\big{(}Wu_{0}+u_{1}\big{)}\end{bmatrix}=\begin{bmatrix}e_{j}^{\top}&0\\ e_{j}^{\top}W&e_{j}\end{bmatrix}\begin{bmatrix}u_{0}\\ u_{1}\end{bmatrix}.$
The general case follows from the same reasoning, similarly as in the proof of Thm. 3.
Appendix E Results for Time-Varying Graphs
We see from the proofs that all the results can be straightforwardly generalized to time-varying gossip matrices $W_{t}$, $t\in[T]$. For example, in case of the non-adaptive compositions with secure summation, we would have Theorem 1 with
$\widetilde{x}_{T}=\begin{bmatrix}e_{i}^{\top}W_{0}e_{j}\\ e_{i}^{\top}(W_{1}\cdot W_{0}+W_{0})e_{j}\\ \vdots\\ e_{i}^{\top}\left(\prod\limits_{i=0}^{T-1}W_{i}+e_{i}^{\top}\prod\limits_{i=0}^{T-2}W_{i}+\cdots+W_{0}\right)e_{j}\end{bmatrix}$
and
$H_{T}:=\begin{bmatrix}e_{i}^{\top}W_{0}&0&0&\dots&0\\ e_{i}^{\top}W_{0}\cdot W_{1}&e_{i}^{\top}W_{1}&0&\dots&0\\ e_{i}^{\top}W_{0}\cdot W_{1}\cdot W_{2}&e_{i}^{\top}W_{1}\cdot W_{2}&e_{i}^{\top}W_{2}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{i}^{\top}\prod\limits_{i=0}^{T-1}W_{i}&e_{i}^{\top}\prod\limits_{i=0}^{T-2}W_{i}&\dots&e_{i}^{\top}W_{T-2}\cdot W_{T-1}&e_{i}^{\top}W_{T-1}\end{bmatrix}$
The vector $\widetilde{x}_{T}$ and the matrix $H_{T}$ of Theorems 3 and 4 can be modified analogously.
Appendix F Details for Experiments
F.1 Additional Details
  • For each model, both when using the DP Decentralized Averaging and DP-FedAvg, we optimize the learning rate using the grid $\{10^{-i/2}\}$, $i\in\mathbb{Z}$.
  • In case of the gossip averaging, the test accuracies are evaluated on a model of a randomly chosen node.
  • The experiments on the logistic regression including all the learning rate tuning took around 48 hours using 8 RTX 3080 GPUs.
F.2 Baseline Method for Experiments
As a baseline algorithm, we consider the full batch version of the DP Federated Averaging algorithm depicted in Alg. 11. (Ponomareva et al.,, 2023; McMahan et al.,, 2017)
Algorithm 1 Differentially Private Federated Averaging (DP-FedAvg) with Record-Level DP
1:  Inputs: number of clients $n$, number of rounds $T$, local learning rate $\eta$, clipping norm $C$, noise parameter $\sigma$, initial global model $w^{0}$
2:  for each round $t=0$ to $T-1$ do
3:     for each client $i$ in parallel do
4:        Initialize local model: $w_{i}^{t}\leftarrow w^{t}$
5:        Perform DP gradient descent step on $D_{i}$, with clipping constant $C$, learning rate $\eta$ and noise multiplier $\sigma/\sqrt{n}$ obtaining updated model $w_{i}^{t}$
6:        Compute model update: $\Delta_{i}^{t}\leftarrow w_{i}^{t}-w^{t}$
7:     end for
8:     Server aggregates updates:
$\tilde{\Delta}^{t}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\Delta_{i}^{t}.$
9:     Update global model:
$w^{t+1}\leftarrow w^{t}+\tilde{\Delta}^{t}$
and distribute it to clients.
10:  end for
11:  Output: Final global model $w^{T}$
F.3 Gossip Averaging Algorithm
In the experiments, we consider the full decentralized DP gradient descent shown in Alg. 10.
Algorithm 2 Differentially Private Gossip Averaging (DP-GossipAvg)
1:  Inputs: number of clients $n$, number of rounds $T$, local learning rate $\eta$, clipping norm $C$, noise multiplier $\sigma$, gossip matrix $W\in\mathbb{R}^{n\times n}$, initial local models $\{w_{i}^{0}\}_{i=1}^{n}$
2:  for each round $t=0$ to $T-1$ do
3:     for each client $i$ in parallel do
4:        Perform local DP gradient descent on $D_{i}$ with clipping constant $C$, learning rate $\eta$ and noise multiplier $\sigma$ to obtain updated local model $w_{i}^{t}$
5:     end for
6:     for each client $i$ in parallel do
7:        Gossip averaging:
$w_{i}^{t+1}\leftarrow\sum_{j=1}^{n}W_{ij}w_{j}^{t}$
8:     end for
9:  end for
10:  Output: Final local models $\{w_{i}^{T}\}_{i=1}^{n}$

### v2

@@ Appendix A Proof of Lemma 6
[…]
@@
[…]
@@ Appendix B Proof of Lemma 7
[…]
where $\widehat{u}\sim\mathcal{N}(0,\sigma^{2}I_{T})$. ∎
Appendix C Proof of Thm. 8
@@
For an upper bound of the sensitivity, we need to find the maximum of $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ when $\Delta x$ is in the convex and compact domain
$\mathcal{D}_{j}:=\{\alpha\otimes e_{j}:\alpha\in[-1,1]^{T}\}$
We see that maximizing $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ is equivalent to finding the maximum of the quadratic form $(\Delta x)^{\top}P\Delta x$ in $\mathcal{D}_{j}$ where the projector matrix $P=V_{r}V_{r}^{\top}$ is positive semi-definite and a projection on to the row space of $H_{T}$. Thus, the optimum is found from somewhere at the boundary of $\mathcal{D}_{j}$, i.e., it holds that $\Delta x=\alpha\otimes e_{j}$, where $\alpha=\{-1,1\}^{T}$.
For the inequality, we use the trivial inequality
$\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}=(\Delta x)^{\top}V_{r}V_{r}^{\top}(\Delta x)\leq\mathbf{1}_{T}^{\top}\left|V_{r}V_{r}^{\top}\right|\mathbf{1}_{T}.$
Appendix D Proof of Lemma 11
@@
Since $W$ is primitive and row-stochastic, $1$ is a simple eigenvalue with right eigenvector $\mathbf{1}$ and left eigenvector $\pi^{\top}$, and all other eigenvalues satisfy $|\lambda|<1$ (Perron-Frobenius). Choose a basis matrix S=[ 1B]S=[\ $\mathbf${1}\ \ B\ ] for some $B\in\mathbb{R}^{n\times n-1}$ so that the first row of $S^{-1}$ is $\pi^{\top}$. Then
$S^{-1}WS\;=\;\begin{bmatrix}1&0\\ 0&T^{\prime}\end{bmatrix}$,
and taking powers gives
$W^{t}\;=\;S\begin{bmatrix}1&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}\;=\;\underbrace{\mathbf{1}\,\pi^{\top}}_{=\ \Pi}\;+\;\underbrace{S\begin{bmatrix}0&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}}_{=:R^{t}}.$
To bound $\|R^{t}\|$, fix any matrix norm $\|\cdot\|$. Since $\rho(T^{\prime}):=\max\{|\lambda|:\lambda\in\mathrm{spec}(T^{\prime})\}<1$, the Gelfand formula implies that for any $\rho$ with $\rho(T^{\prime})<\rho<1$ there exists $C<\infty$ such that $\|(T^{\prime})^{t}\|\leq C\,\rho^{t}$ for all $t\geq 1$. Therefore
$\|R^{t}\|\;=\;\left\|S\begin{bmatrix}0&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}\right\|\;\leq\;\|S\|\,\|S^{-1}\|\,\|(T^{\prime})^{t}\|\;\leq\;\big(\|S\|\,\|S^{-1}\|\;C\big)\,\rho^{t}.$
Setting $M:=\|S\|\,\|S^{-1}\|\,C$ gives the claim. ∎
Appendix E Proof Theorem 13
Proof.
First, from the compact SVD we see that for any matrix $A$, the product $A^{+}A$ is the orthogonal projector onto $\mathrm{row}(A)$. Set $P_{T}:=H_{T}^{+}H_{T}$ to be the orthogonal projection onto row space of $H_{T}$. Let $c\in\{\pm 1\}^{T}$ and define $z:=c\otimes e_{j}\in\mathbb{R}^{nT}$. Our strategy is to study the sensitivity
$\Delta_{T}(c,j\to i)=\|P_{T}z\|_{2}$
and derive a bound for it that is independent of $c$.
Because $W$ is primitive and row-stochastic, by Lemma 11 there exist constants $M>0$ and $\rho\in(0,1)$ such that
$W^{t}=\Pi+R^{t},\qquad\Pi:=\mathbf{1}\pi^{\top},\qquad\|R^{t}\|_{2}\leq M\rho^{t}\quad\forall t\geq 0.$ (E.1)
The $(t,s)$ block of $H_{T}$ (with $0\leq s\leq t\leq T-1$) is given by
$H_{T}(t,s)=e_{i}^{\top}W^{t-s}=e_{i}^{\top}\Pi+e_{i}^{\top}R^{t-s}=\pi^{\top}+e_{i}^{\top}R^{t-s}.$
Define the stationary part of $H_{T}$ as $\bar{H}_{T}$ such that blocks of $\bar{H}_{T}$ are given as
$\bar{H}_{T}(t,s)=\pi^{\top}\quad\text{ for }s\leq t,\qquad\bar{H}_{T}(t,s)=0\quad\text{for }s>t.$
Furthermore, define the transient part
$K_{T}:=H_{T}-\bar{H}_{T}.$
Thus
$H_{T}=\bar{H}_{T}+K_{T}.$
We clearly have that
$\mathrm{row}(\bar{H}_{T})=\{(\alpha_{0}\pi,\alpha_{1}\pi,\dots,\alpha_{T-1}\pi)\in\mathbb{R}^{nT}:\alpha_{t}\in\mathbb{R}\}$
and we can take as an orthonormal basis for $\mathrm{row}(\bar{H}_{T})$ for example $\{e_{1}\otimes\pi/\|\pi\|_{2},\ldots,e_{T}\otimes\pi/\|\pi\|_{2}\}.$ Therefore the orthogonal projection onto $\mathrm{row}(\bar{H}_{T})$ is given by
$\bar{P}_{T}z=\sum_{\ell=1}^{T}\langle e_{\ell}\otimes\pi/\|\pi\|_{2},c\otimes e_{j}\rangle=c\;\otimes\;\frac{\pi_{j}}{\|\pi\|_{2}^{2}}\pi.$
and we have that
$\|\bar{P}_{T}z\|_{2}^{2}=\sum_{t=0}^{T-1}\left\|c_{t}\,\frac{\pi_{j}}{\|\pi\|_{2}^{2}}\pi\right\|_{2}^{2}=T\,\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$
We next study the transient part and show using standard perturbation theory for projection matrices that the distance between the projectors $P_{T}$ and $\bar{P}_{T}$ satisfies
$\|P_{T}-\bar{P}_{T}\|_{2}=O(1/\sqrt{T}).$
This rate is derived using the Wedin theorem which states that the distance between the row spaces $\mathrm{row}(H_{T})$ and $\mathrm{row}(\bar{H}_{T})$ is a function of $\|K_{T}\|_{2}$ (since the perturbation $H_{T}-\bar{H}_{T}=K_{T}$) and the smallest non-zero singular value of $\bar{H}_{T}$, $\sigma_{\min}(\bar{H}_{T})$.
From the Toeplitz structure it follows that clearly $\|K_{T}\|_{2}$ equals the 2-norm of the first column block of $K_{T}$. Moreover, from the transient behavior stated in Eq. (E.1) and a geometric sum it follows that $\|K_{T}\|_{2}=O(1)$. We see that
$\bar{H}_{T}=U\otimes\pi^{\top}$
for
$U=\begin{pmatrix}1&0&\cdots&0\\ 1&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\cdots&1\end{pmatrix}$.
The singular values of a Kronecker product $A\otimes B$ are the products of the singular values of $A$ and $B$. Thus, the non-zero singular values of $\bar{H}_{T}$ are:
$\sigma_{k}(\bar{H}_{T})=\sigma_{m}(U)\cdot\sigma_{l}(\pi^{\top}).$
Since $\pi^{\top}$ has rank $1$, $\sigma_{1}(\pi^{\top})=\|\pi^{\top}\|_{2}=\|\pi\|_{2}.$ We also know that the smallest singular value of $U$ is given by
$\sigma_{\min}(U)=\sqrt{T}.$
Thus, $\sigma_{\min}(\bar{H}_{T})=\sqrt{T}\|\pi\|_{2}$, and by the Wedin theorem we have that
$\|P_{T}-\bar{P}_{T}\|_{2}\leq\frac{2\|K_{T}\|_{2}}{\sigma_{\min}(\bar{H}_{T})}=O(1/\sqrt{T}).$
Let $\mathcal{E}_{T}=P_{T}-\bar{P}_{T}$, so $\|\mathcal{E}_{T}\|_{2}=O(1/\sqrt{T})$.
We expand $\|P_{T}z\|_{2}^{2}$ using $P_{T}=\bar{P}_{T}+\mathcal{E}_{T}$. Recall $z=c\otimes e_{j}$ and $\|z\|_{2}=\sqrt{T}$:
$\|P_{T}z\|_{2}^{2}=\|(\bar{P}_{T}+\mathcal{E}_{T})z\|_{2}^{2}=\|\bar{P}_{T}z\|_{2}^{2}+2(\bar{P}_{T}z)^{\top}(\mathcal{E}_{T}z)+\|\mathcal{E}_{T}z\|_{2}^{2}.$ (E.2)
Using a norm bound on $\mathcal{E}_{T}$, we have
$|2(\bar{P}_{T}z)^{\top}(\mathcal{E}_{T}z)|$ $\leq 2\|\bar{P}_{T}z\|_{2}\|\mathcal{E}_{T}z\|_{2}$
$\leq 2\|\bar{P}_{T}z\|_{2}\|\mathcal{E}_{T}\|_{2}\|z\|_{2}$
$\leq 2\left(\sqrt{T}\frac{\pi_{j}}{\|\pi\|_{2}}\right)\cdot O(1/\sqrt{T})\cdot\sqrt{T}=O(\sqrt{T}).$
Furthermore,
$\|\mathcal{E}_{T}z\|_{2}^{2}\leq\|\mathcal{E}_{T}\|_{2}^{2}\|z\|_{2}^{2}=O(1/T)\cdot T=O(1).$
Substituting these bounds back into Eq. (E.2):
$\|P_{T}z\|_{2}^{2}=T\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(\sqrt{T}).$
This holds for any choice of $c\in\{\pm 1\}^{T}$. The maximum is thus:
$(\Delta_{T}(j\to i))^{2}=\max_{c\in\{\pm 1\}^{T}}\|P_{T}z\|_{2}^{2}=T\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(\sqrt{T}).$
Dividing by $T$ and taking the limit:
$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\lim_{T\to\infty}\left(\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(1/\sqrt{T})\right)=\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$